OpenStudy (anonymous):

Simplify the following (these are examples to help me work out the actual problems). ln1 lne^7 log(subscript3)(1/3)

OpenStudy (anonymous):

first off it is always true that \(b^0=1\) and therefore \(\log_b(1)=0\)

OpenStudy (anonymous):

another way to put this is since \(e^0=1\) we know \(\ln(1)=0\) since \(\ln(x)=\log_e(x)\)

OpenStudy (anonymous):

since \(e^x\) and \(\ln(x)\) are inverse functions, the composite is always the identity in other words it is always true that \[e^{\ln(x)}=x\]and \[ ln(e^x)=x\]

OpenStudy (anonymous):

since \(b^{-1}=\frac{1}{b}\) we know \[\log_b(\frac{1}{b})=-1\]

OpenStudy (anonymous):

I need a little further elabortaion on the last one, how would the variable still work in the equation? would it become negative 3 due to b^-1. Eh, if you could workout a different example for me, that would be great.